OpenStudy (anonymous):

Let f(x) = x2 - 8x + 5. Find f(-1).

OpenStudy (zale101):

replce -1 with x and solve the equation

OpenStudy (zale101):

replace

OpenStudy (zale101):

(-1)^2 - 8(-1) + 5.

jimthompson5910 (jim_thompson5910):

Like Zale is saying: replace every copy of 'x' with -1 and then follow PEMDAS to evaluate and simplify. \[\large f(x) = x^2 - 8x + 5\] \[\large f(-1) = (-1)^2 - 8(-1) + 5\] \[\large f(-1) = 1 - 8(-1) + 5\] \[\large f(-1) = 1 + 8 + 5\] \[\large f(-1) = 9 + 5\] \[\large f(-1) = 14\]

OpenStudy (anonymous):

@jim_thompson5910 thankyou !

jimthompson5910 (jim_thompson5910):

you're welcome